OpenStudy (anonymous):

find each function value, if it exists. f(t)=√(t^2+1)

jhonyy9 (jhonyy9):

sqrt all this (t2 +1) ???

jhonyy9 (jhonyy9):

how is correct ?

OpenStudy (anonymous):

huh?

jhonyy9 (jhonyy9):

write this newly please

jhonyy9 (jhonyy9):

f(t)=sqrt(t2 +1)

jhonyy9 (jhonyy9):

right ?

jhonyy9 (jhonyy9):

so than sqrt(t2 +1)=0 than squared both sides will get t2+1=0 and so t2= -1 so t =- i